# SQL-R1: Training Natural Language to SQL Reasoning Model By Reinforcement Learning

### v3

@@ 1 Introduction
Natural Language to SQL (NL2SQL, or Text2SQL) converts natural language questions (NL) into structured SQL statements, simplifying database interaction without requiring database expertise (DBLP:journals/corr/abs-2408-05109, ; hong2024next, ). Recent advancements in NL2SQL have significantly enhanced the level of human-computer interaction within database query applications and contribute to a wide range of data science analysis tasks (DBLP:journals/debu/Luo00LZY20, ; DBLP:journals/tkde/LuoQCTLL22, ). Current NL2SQL models mainly focus on optimizing workflows and their components, such as schema linking (resdsql, ; cao2024rsl, ), content retrieval (chess, ), generation correction (dinsql, ; macsql, ; c3sql, ; ma2024plug, ; liu2025nl2sql, ).
Despite these advancements, improving the NL2SQL system’s inference performance in complex database scenarios remains a considerable challenge. As shown in Figure 1, schema complexity may lead to generation errors in processing multi-table joins and nested queries, and it is difficult for individually trained models to think and process complex semantics. Currently, a significant portion of NL2SQL research is devoted to training open-source large language models (LLMs) by supervised fine-tuning (SFT) (dtssql, ; codes, ; catsql, ) to achieve accuracy at a smaller model scale compared to approaches using closed-source LLMs (e.g., GPT-4, GPT-4o) (dailsql, ; dinsql, ; c3sql, ). However, SFT relies on the database schema structure and the training data scale. This may lead to the existing model’s instability in domain adaptation and generalization in new database environments. Additionally, the lack of interpretability of NL2SQL reasoning logic limits the application of the model in high-risk fields, such as finance and healthcare.
@@
Recently, reinforcement learning (RL) has shown great potential in training the reasoning ability of LLMs in recent research. Compared with supervised fine-tuning, reinforcement learning can dynamically adjust the decision-making strategy of the LLMs through interaction with the environment, thereby achieving superior performance in complex reasoning tasks (deepseekr1, ). RL-based methods have proven effective in enhancing model reasoning and generalization capabilities in financial reasoning (finr1, ), search engines (searchr1, ), and mathematical reasoning (logicrl, ; openr1, ).
@@ 2.2.1 Source
Currently, we utilize the SynSQL-2.5M (omnisql, ) dataset as primary data source, which is the first million-scale synthetic NL2SQL dataset, encompassing over 2.5 million diverse and high-quality data samples. Each sample consists of a quadruple comprising a database, a natural language question, an SQL query, and a chain-of-thought (CoT) solution. The dataset features more than 16,000 synthetic databases across various domains, thereby ensuring extensive coverage of realistic scenarios. SynSQL-2.5M includes a wide range of SQL complexity levels, from simple single-table queries to intricate multi-table joins, functions, and common table expressions.
@@ 2.3.1 Supervised Fine-Tuning
In this study, we conduct SFT on the Qwen2.5-Coder-7B-Instruct model to enhance the model’s capacity for instruction adherence and generation within the NL2SQL domain. We investigate two distinct strategies for SFT cold start training. The first one employs raw instructions focusing exclusively on SQL generation. We leverage the existing OmniSQL-7B (omnisql, ) checkpoint for the reference. The second strategy utilizes full fine-tuning and reasoning generation instructions promoting the development of compliant thought processes alongside final answers.
@@ 2.3.2 Reinforcement Training
In the reinforcement learning phase, we employ the Group Relative Policy Optimization (GRPO) algorithm to enhance our training protocol, which obviates the need for the value model, operates with less memory requirements, and facilitates a clear definition of reward targets, rendering it an optimal choice for the effective optimization of the NL2SQL policy model (grpo, ).
@@ 2.3.3 Reward Function Design
$S_{e}=\begin{cases}2,&\text{if SQL candidate is executable}\\ 0,&\text{if format is incorrect}\\ -2,&\text{if SQL candidate is not executable}\end{cases}$ (3)
@@
$S_{r}=\begin{cases}3,&\text{if query result is correct}\\ 0,&\text{if format is incorrect or SQL candidate is not executable}\\ -3,&\text{if query result is incorrect}\end{cases}$ (4)
@@
$S_{l}=\begin{cases}0.5\times S_{tl}+S_{al},&\text{if query result is correct and $len_{response}<=$ {MAX LENGTH}}\\ 0.5+S_{al},&\text{if query result is correct and $len_{response}>$ {MAX LENGTH}}\\ 0,&\text{other cases}\\ \end{cases}$ (5)
@@ 3.1 Setup
We evaluated the proposed SQL-R1 and related NL2SQL models on two benchmarks, Spider (spider, ) and BIRD (bird, ). Spider comprises 10,181 questions paired with 5,693 complex SQL queries from 200 databases and 138 domains. BIRD comprises 12,751 NL2SQL pairs encompassing 95 databases from 37 specialized domains.
@@
Currently, SQL-R1 is mainly built on Qwen2.5-Coder series models (qwen2.5, ). For the SFT training, we set the learning rate as 5e-5; batch size as 1. For the RL training, we set the learning rate as 3e-7, rollout of actor model as 8; max response length as 2048. For inference, we set the count of SQL candidates as 8 and the temperature as 0.8.
For all NL2SQL data samples in the dataset, we first convert them into suitable input-output sequence pairs for training. Specifically, the input sequence includes natural language questions and related database schemas. Inspired by previous research (chess, ; omnisql, ; yang2024synthesizing, ; rajkumar2022evaluating, ), the database schema was formatted as a CREATE TABLE statement with supplementary annotations including column attribute descriptions and representative values. Currently, annotations of representative values will not be added during the training phase for the time being to enhance the model’s exploration ability during the reinforcement learning phase.
@@ 3.2 Main Results
CodeS (codes, ) CodeS-15B - 84.9 79.4 57.0
DTS-SQL (dtssql, ) Deepseek-Coder-7B - 85.5 84.4 55.8
CHESS (chess, ) Deepseek-Coder-33B - - 87.2 61.5
Alpha-SQL (alphasql, ) Qwen2.5-Coder-7B Self-Consistency 84.0 - 66.8
SQL-o1 (sqlo1, ) Qwen2.5-Coder-7B Self-Consistency 84.7 85.1 66.7
OmniSQL (omnisql, ) Qwen2.5-Coder-7B Self-Consistency 85.5 88.9 66.1
DeepRetrieval (deepretrieval, ) Qwen2.5-Coder-7B - - 76.1 56.0
Reasoning-SQL (reasoningsql, ) Qwen2.5-Coder-14B Self-Consistency 81.4 - 65.3
C3-SQL (c3sql, ) GPT-3.5-Turbo Self-Consistency 82.0 82.3 -
DIN-SQL (dinsql, ) GPT-4 - 82.8 85.3 -
DAIL-SQL (dailsql, ) GPT-4 Self-Consistency 83.6 86.2 54.8
MAC-SQL (macsql, ) GPT-4 Self-Consistency 86.8 82.8 59.4
SuperSQL (supersql, ) GPT-4 Self-Consistency 84.0 87.0 58.5
MCTS-SQL (mctssql, ) GPT-4o - 88.7 86.6 69.4
OpenSearch-SQL (opensearchsql, ) GPT-4o Self-Consistency - 87.1 69.3
CHASE-SQL (chasesql, ) Gemimi-1.5-Pro - - 87.6 73.0
@@
The results presented in Table 1 underscore the performance of SQL-R1 across various foundational models. Specifically, when trained on the Qwen2.5-Coder-3B, SQL-R1 achieved an execution accuracy of 78.1% on the Spider development dataset, 78.9% on the Spider test dataset, and 54.6% on the BIRD development dataset, demonstrating robust performance even with a smaller base model. Moreover, SQL-R1 exhibits a marked enhancement in accuracy when utilized with larger models; for instance, the Qwen2.5-Coder-7B model yielded accuracies of 87.6%, 88.7%, and 63.1% across the same Spider development, Spider test, and BIRD development datasets, respectively. The deployment of the even larger Qwen2.5-Coder-14B model further affirmed SQL-R1’s capabilities, achieving high accuracy rates of 86.7%, 88.1%, and 67.1% on these datasets. In comparison to other solutions leveraging closed-source models, such as GPT-4 and GPT-4o, SQL-R1 demonstrates competitive and often superior performance. Notably, In comparison to contemporaneous exploration of reasoning-based NL2SQL (e.g., DeepRetrieval (deepretrieval, ) and Reasoning-SQL (reasoningsql, )), SQL-R1 surpasses them on the Spider-Dev and BIRD-Dev datasets. This performance highlights the efficacy of our proposed approach in utilizing diverse base models to attain state-of-the-art results on complex NL2SQL reasoning tasks. Additional experimental results on other NL2SQL benchmarks can be found in Section B.1 and Table 4.
@@
As illustrated in Figure 2, we investigated the relationship between performance and model size utilizing on BIRD development dataset. The comparative analysis encompasses various model types, including NL2SQL models, reasoning LLMs and general LLMs. For GPT-4, GPT-4o, GPT-4o-mini and Gemni-1.5-Pro, we refer to the parametric description of (abacha2024medec, ; omnisql, ; alphasql, ). The findings demonstrate that when utilizing the 7B model as a foundational model, SQL-R1 attains accuracy levels that surpass those of larger-scale models, particularly. This underscores the efficacy of the proposed model in optimizing the performance of NL2SQL while ensuring that cost efficiency is maintained. Additionally, Section B.2 provides some detailed results of different base models.
@@ 3.4 Ablation Study of Reward Components
We conducted an ablation experiment on the BIRD development dataset to verify the proposed reinforcement learning reward mechanism’s effectiveness. The experiment sequentially removed individual reward components from the comprehensive reward function while maintaining the parameter settings established in Section 3.1. As presented in Table 3, the result demonstrates that omitting any reward component from the original reward function adversely impacts inference performance. This underscores the critical importance of execution feedback and result reward in the model training process. Section B.3 and B.4 provide more details of the ablation study.
@@ 4 Related Works
Recent advancements in large language models have significantly propelled the development of NL2SQL translation techniques (DBLP:journals/corr/abs-2408-05109, ; katsogiannis2023survey, ; hong2024next, ). Current research focuses on optimizing various components of the NL2SQL workflow, including pre-processing modules like schema linking (resdsql, ), translation strategies (scprompt, ; zeronl2sql, ; natsql, ; wolfson2020break, ; eyal2023semantic, ; gao2022towards, ; petsql, ), and post-processing modules like SQL correction (liu2025nl2sql, ; dinsql, ; macsql, ) and self-consistency for SQL selection (dailsql, ; c3sql, ; omnisql, ). However, existing NL2SQL models primarily rely on supervised fine-tuning, which may exhibit instability in domain adaptation and new scenario fitting, as they heavily depend on the knowledge of database schema and structure from the training data (dtssql, ; catsql, ; ratsql, ). Additionally, the lack of interpretability in the reasoning logic of these NL2SQL models limits their application in high-risk scenarios or complex database requirements (spider2, ). To address these limitations, we explore the application of reinforcement learning to enhance the reasoning and generalization capabilities of NL2SQL models. By dynamically adjusting the model’s decision-making strategy through interaction with the environment, we aim to improve its performance in complex database scenarios and ensure better domain adaptation and interpretability.
@@
Recent research on reinforcement learning for LLMs has increasingly concentrated on enhancing reasoning capabilities and optimizing interaction with the external environment to improve proficiency in complex, multi-step reasoning tasks (wei2022chain, ; plaat2024reasoning, ). In recent months, DeepSeek-R1 (deepseekr1, ) have illustrated the effectiveness of group relative policy optimization (GRPO), a novel and efficient reinforcement algorithm in cultivating robust reasoning behaviors in LLMs. By focusing exclusively on the correctness of the final answer, GRPO facilitates the internalization of intermediate reasoning steps without the necessity of explicit supervision, achieving notable success across various domains, including mathematics (shao2024deepseekmath, ), finance (finr1, ), programming (pan2025metaspatial, ), user interface (uir1, ; guir1, ) and vision tasks (visionr1, ). Concurrently, some studies have explored integrating external tools, such as retrieval (deepretrieval, ) and search engines (searchr1, ), into reinforcement learning frameworks for LLM reasoning, primarily focusing on training models to effectively employ a single, general-purpose tool to support specific reasoning chains. In light of these advancements, we aim to extend LLM reasoning capabilities to the NL2SQL task by adopting a reinforcement learning approach that dynamically adjusts the model’s decision-making strategy through interaction with the environment, thereby enhancing its ability to manage complex semantics and improve generalization and domain adaptation capabilities.
@@ B.2 Analysis and Comparison of Base LLMs
B.3 Ablation Study of SQL Candidate Rollouts in Reasoning

### v4

@@ 1 Introduction
Natural Language to SQL (NL2SQL, or Text2SQL) converts natural language questions (NL) into structured SQL statements, simplifying database interaction without requiring database expertise [1, 2]. Recent advancements in NL2SQL have significantly enhanced the level of human-computer interaction within database query applications and contribute to a wide range of data science analysis tasks [3, 4]. Current NL2SQL models mainly focus on optimizing workflows and their components, such as schema linking [5, 6], content retrieval [7], generation correction [8, 9, 10, 11, 12].
Despite these advancements, improving the NL2SQL system’s inference performance in complex database scenarios remains a considerable challenge. As shown in Figure 1, schema complexity may lead to generation errors in processing multi-table joins and nested queries, and it is difficult for individually trained models to think and process complex semantics. Currently, a significant portion of NL2SQL research is devoted to training open-source large language models (LLMs) by supervised fine-tuning (SFT) [13, 14, 15] to achieve accuracy at a smaller model scale compared to approaches using closed-source LLMs (e.g., GPT-4, GPT-4o) [16, 8, 10]. However, SFT relies on the database schema structure and the training data scale. This may lead to the existing model’s instability in domain adaptation and generalization in new database environments. Additionally, the lack of interpretability of NL2SQL reasoning logic limits the application of the model in high-risk fields, such as finance and healthcare.
@@
Recently, reinforcement learning (RL) has shown great potential in training the reasoning ability of LLMs in recent research. Compared with supervised fine-tuning, reinforcement learning can dynamically adjust the decision-making strategy of the LLMs through interaction with the environment, thereby achieving superior performance in complex reasoning tasks [17]. RL-based methods have proven effective in enhancing model reasoning and generalization capabilities in financial reasoning [18], search engines [19], and mathematical reasoning [20, 21].
@@ 2.2.1 Source
Currently, we utilize the SynSQL-2.5M [22] dataset as primary data source, which is the first million-scale synthetic NL2SQL dataset, encompassing over 2.5 million diverse and high-quality data samples. Each sample consists of a quadruple comprising a database, a natural language question, an SQL query, and a chain-of-thought (CoT) solution. The dataset features more than 16,000 synthetic databases across various domains, thereby ensuring extensive coverage of realistic scenarios. SynSQL-2.5M includes a wide range of SQL complexity levels, from simple single-table queries to intricate multi-table joins, functions, and common table expressions.
@@ 2.3.1 Supervised Fine-Tuning
In this study, we conduct SFT on the Qwen2.5-Coder-7B-Instruct model to enhance the model’s capacity for instruction adherence and generation within the NL2SQL domain. We investigate two distinct strategies for SFT cold start training. The first one employs raw instructions focusing exclusively on SQL generation. We leverage the existing OmniSQL-7B [22] checkpoint for the reference. The second strategy utilizes full fine-tuning and reasoning generation instructions promoting the development of compliant thought processes alongside final answers.
@@ 2.3.2 Reinforcement Training
In the reinforcement learning phase, we employ the Group Relative Policy Optimization (GRPO) algorithm to enhance our training protocol, which obviates the need for the value model, operates with less memory requirements, and facilitates a clear definition of reward targets, rendering it an optimal choice for the effective optimization of the NL2SQL policy model [23].
@@ 2.3.3 Reward Function Design
[…]
@@
[…]
@@
[…]
@@ 3.1 Setup
We evaluated the proposed SQL-R1 and related NL2SQL models on two benchmarks, Spider [24] and BIRD [25]. Spider comprises 10,181 questions paired with 5,693 complex SQL queries from 200 databases and 138 domains. BIRD comprises 12,751 NL2SQL pairs encompassing 95 databases from 37 specialized domains.
@@
Currently, SQL-R1 is mainly built on Qwen2.5-Coder series models [26]. For the SFT training, we set the learning rate as 5e-5; batch size as 1. For the RL training, we set the learning rate as 3e-7, rollout of actor model as 8; max response length as 2048. For inference, we set the count of SQL candidates as 8 and the temperature as 0.8.
For all NL2SQL data samples in the dataset, we first convert them into suitable input-output sequence pairs for training. Specifically, the input sequence includes natural language questions and related database schemas. Inspired by previous research [7, 22, 27, 28], the database schema was formatted as a CREATE TABLE statement with supplementary annotations including column attribute descriptions and representative values. Currently, annotations of representative values will not be added during the training phase for the time being to enhance the model’s exploration ability during the reinforcement learning phase.
@@ 3.2 Main Results
CodeS [14] CodeS-15B - 84.9 79.4 57.0
DTS-SQL [13] Deepseek-Coder-7B - 85.5 84.4 55.8
CHESS [7] Deepseek-Coder-33B - - 87.2 61.5
Alpha-SQL [29] Qwen2.5-Coder-7B Self-Consistency 84.0 - 66.8
SQL-o1 [30] Qwen2.5-Coder-7B Self-Consistency 84.7 85.1 66.7
OmniSQL [22] Qwen2.5-Coder-7B Self-Consistency 85.5 88.9 66.1
DeepRetrieval [31] Qwen2.5-Coder-7B - - 76.1 56.0
Reasoning-SQL [32] Qwen2.5-Coder-14B Self-Consistency 81.4 - 65.3
C3-SQL [10] GPT-3.5-Turbo Self-Consistency 82.0 82.3 -
DIN-SQL [8] GPT-4 - 82.8 85.3 -
DAIL-SQL [16] GPT-4 Self-Consistency 83.6 86.2 54.8
MAC-SQL [9] GPT-4 Self-Consistency 86.8 82.8 59.4
SuperSQL [33] GPT-4 Self-Consistency 84.0 87.0 58.5
MCTS-SQL [34] GPT-4o - 88.7 86.6 69.4
OpenSearch-SQL [35] GPT-4o Self-Consistency - 87.1 69.3
CHASE-SQL [36] Gemimi-1.5-Pro - - 87.6 73.0
@@
The results presented in Table 1 underscore the performance of SQL-R1 across various foundational models. Specifically, when trained on the Qwen2.5-Coder-3B, SQL-R1 achieved an execution accuracy of 78.1% on the Spider development dataset, 78.9% on the Spider test dataset, and 54.6% on the BIRD development dataset, demonstrating robust performance even with a smaller base model. Moreover, SQL-R1 exhibits a marked enhancement in accuracy when utilized with larger models; for instance, the Qwen2.5-Coder-7B model yielded accuracies of 87.6%, 88.7%, and 63.1% across the same Spider development, Spider test, and BIRD development datasets, respectively. The deployment of the even larger Qwen2.5-Coder-14B model further affirmed SQL-R1’s capabilities, achieving high accuracy rates of 86.7%, 88.1%, and 67.1% on these datasets. In comparison to other solutions leveraging closed-source models, such as GPT-4 and GPT-4o, SQL-R1 demonstrates competitive and often superior performance. Notably, In comparison to contemporaneous exploration of reasoning-based NL2SQL (e.g., DeepRetrieval [31] and Reasoning-SQL [32]), SQL-R1 surpasses them on the Spider-Dev and BIRD-Dev datasets. This performance highlights the efficacy of our proposed approach in utilizing diverse base models to attain state-of-the-art results on complex NL2SQL reasoning tasks. Additional experimental results on other NL2SQL benchmarks (e.g., Spider2.0) can be found in Table 4, Table 5 and Section B.1.
@@
As illustrated in Figure 2, we investigated the relationship between performance and model size utilizing on BIRD development dataset. The comparative analysis encompasses various model types, including NL2SQL models, reasoning LLMs and general LLMs. For GPT-4, GPT-4o, GPT-4o-mini and Gemni-1.5-Pro, we refer to the parametric description of [37, 22, 29]. The findings demonstrate that when utilizing the 7B model as a foundational model, SQL-R1 attains accuracy levels that surpass those of larger-scale models, particularly. This underscores the efficacy of the proposed model in optimizing the performance of NL2SQL while ensuring that cost efficiency is maintained. Additionally, Section B.2 provides some detailed results of different base models.
@@ 3.4 Ablation Study of Reward Components
We conducted an ablation experiment on the BIRD development dataset to verify the proposed reinforcement learning reward mechanism’s effectiveness. The experiment sequentially removed individual reward components from the comprehensive reward function while maintaining the parameter settings established in Section 3.1. As presented in Table 4, the result demonstrates that omitting any reward component from the original reward function adversely impacts inference performance. This underscores the critical importance of execution feedback and result reward in the model training process. Section B.4 and B.5 provide more details of the ablation study.
@@
Table 4: Execution accuracy (%) of different models on Spider2.0 SQLite subset.
Model Accuracy (%)
Qwen2.5-Coder-7B 2.2
GPT-4o 15.6
Deepseek-V3 15.6
OmniSQL-7B 10.4
SQL-R1-7B 20.0
@@ 4 Related Works
Recent advancements in large language models have significantly propelled the development of NL2SQL translation techniques [1, 38, 2]. Current research focuses on optimizing various components of the NL2SQL workflow, including pre-processing modules like schema linking [5], translation strategies [39, 40, 41, 42, 43, 44, 45], and post-processing modules like SQL correction [12, 8, 9] and self-consistency for SQL selection [16, 10, 22]. However, existing NL2SQL models primarily rely on supervised fine-tuning, which may exhibit instability in domain adaptation and new scenario fitting, as they heavily depend on the knowledge of database schema and structure from the training data [13, 15, 46]. Additionally, the lack of interpretability in the reasoning logic of these NL2SQL models limits their application in high-risk scenarios or complex database requirements [47]. To address these limitations, we explore the application of reinforcement learning to enhance the reasoning and generalization capabilities of NL2SQL models. By dynamically adjusting the model’s decision-making strategy through interaction with the environment, we aim to improve its performance in complex database scenarios and ensure better domain adaptation and interpretability.
@@
Recent research on reinforcement learning for LLMs has increasingly concentrated on enhancing reasoning capabilities and optimizing interaction with the external environment to improve proficiency in complex, multi-step reasoning tasks [48, 49]. In recent months, DeepSeek-R1 [17] have illustrated the effectiveness of group relative policy optimization (GRPO), a novel and efficient reinforcement algorithm in cultivating robust reasoning behaviors in LLMs. By focusing exclusively on the correctness of the final answer, GRPO facilitates the internalization of intermediate reasoning steps without the necessity of explicit supervision, achieving notable success across various domains, including mathematics [50], finance [18], programming [51], user interface [52, 53] and vision tasks [54]. Concurrently, some studies have explored integrating external tools, such as retrieval [31] and search engines [19], into reinforcement learning frameworks for LLM reasoning, primarily focusing on training models to effectively employ a single, general-purpose tool to support specific reasoning chains. In light of these advancements, we aim to extend LLM reasoning capabilities to the NL2SQL task by adopting a reinforcement learning approach that dynamically adjusts the model’s decision-making strategy through interaction with the environment, thereby enhancing its ability to manage complex semantics and improve generalization and domain adaptation capabilities.
@@ Appendix B Additional Results and Analysis
Table 6: Execution accuracy (%) of different complexity levels on BIRD-Dev dataset.
NL2SQL Method Base Model Simple Moderate Challenging All
CodeS Codes-15B 65.8 48.8 42.4 58.5
DAIL-SQL GPT-4 63.0 45.6 43.1 55.9
SuperSQL GPT-4 66.9 46.5 43.8 58.5
SQL-R1 (Ours) Qwen2.5-Coder-7B 72.1 60.8 51.0 66.6
SQL-R1 (Ours) Qwen2.5-Coder-14B 72.4 59.7 56.5 67.1
B.1 Additional Analysis of Performance on NL2SQL Benchmarks
Performance on other NL2SQL Benchmarks.
Complexity Analysis.
We performed a statistical analysis of accuracy rates across different difficulty levels based on the BIRD-Dev dataset. The results in Table 6 demonstrate that SQL-R1 significantly outperforms the baseline across all three subsets: Simple, Moderate, and Challenging, indicating its robust capability in reasoning and generation for various SQL generation tasks. When we scaled SQL-R1 from 7B to 14B parameters, the accuracy rates for Simple and Moderate levels experienced minimal changes of just 0.3% and a decrease of 1.1%, respectively. However, a noteworthy improvement was observed in the Challenging category, where accuracy increased from 51.0% to 56.5%, marking a gain of 5.5 percentage points. This finding suggests that enhancing model capacity primarily addresses the generalization challenges associated with complex queries. Moreover, it aligns with existing research that indicates larger models are better at capturing long-range dependencies, particularly in scenarios involving multiple table joins, nested subqueries, or complex aggregations, highlighting the importance of SQL-R1’s size in improving performance on sophisticated SQL generation tasks.
@@ B.2 Analysis and Comparison of Base LLMs
B.3 Analysis and Comparison of Efficiency
We assessed the latency of SQL-R1 on the BIRD-Dev dataset, maintaining the same inference settings and evaluation metric as outlined in Section 3.1.
Table 8: Efficiency comparison of different NL2SQL methods on BIRD-dev dataset. The base model of XiYan-SQL and SQL-R1 is Qwen2.5-Coder-7B.
NL2SQL Method
Candidate
Selection
Latency (s) /
Question
Total Tokens (K)
Question
Accuracy (%)
Qwen2.5-Coder-7B Greedy Search 0.3 2.5 58.2
XiYan-SQL Greedy Search 0.5 4.1 62.1
CHESS Greedy Search 251.3 320.8 61.5
SQL-R1 (Ours) Greedy Search 0.4 3.1 63.7
SQL-R1 (Ours) Self-Consistency 1.1 23.1 66.6
As demonstrated in Table 8, the self-consistency method employing 8 candidates adds only an average of 0.7 seconds to each query than greedy search method, while improving execution accuracy by 2.9%, which represents a favorable trade-off. Additionally, We compared SQL-R1 with XiYan-SQL [59] using the same base model and related baselines, and observed that SQL-R1 achieved higher accuracy under the same conditions. In contrast to CHESS [7], which has an execution accuracy of 61.5% and employs a multi-component workflow, the end-to-end SQL-R1 model achieves an accuracy of 66.6%. SQL-R1 not only shows superior execution accuracy but also offers significant advantages in terms of latency and token usage efficiency.
B.4 Ablation Study of SQL Candidate Rollouts in Reasoning